# QZO: A Catalog of 5 Million Quasars from the Zwicky Transient Facility

### v1

@@ 2.4 Other surveys
We compare the ZTF based classification with AllWISE, Gaia EDR3, and PS DR1 by cross-matching the ZTF training datasets in g- and r-bands using a $1\arcsec$ matching distance. From PS, AllWISE, and Gaia, we use griz, W[1-4], and g, bp, rp magnitudes, respectively. We add colors from all magnitude pair combinations within each survey. Additionally, we add parallax and proper motion measurements from Gaia. We do not use error measurements for any of the features, either as inputs or weights for loss functions. Requiring availability of all the cross-matched features, we obtain the following subsets of the ZTF g-band training data: ZTF x PS 98%, ZTF x WISE 77%, ZTF x Gaia 45%. When comparing the classification based on different surveys, we apply the magnitude limits: PS griz 22.0, 21.8, 21.5, 20.9, WISE W[1-2] 17.1, 15.7, Gaia g 21, respectively. Fig. 3 shows a comparison of quasar distributions between the full training sample and the cross-matched subsets. We also cross-match with AllWISE for the inference data, as justified in Sections 4.2 and 4.3. For simplicity, we use WISE to refer to the AllWISE sample.
@@ 3.1 Astromer
We classify the light curves using Astromer (Donoso-Oliva et al., 2023), a deep learning transformer architecture, pre-trained on all the ZTF DR10 g-band data. The model uses attention layers trained in a semi supervised way by removing random windows from the light curves, and then recreating the missing parts. The model is limited to 200 input observation epochs, transforming each observation into a vector of 256 tokens, which results in an encoder output size of $200\times 256$. We fine-tune the model for a classification task by processing the encoder output with other types of layers.
@@ 3.2 Pipeline
We first retrain Astromer using its original training procedure on the g- and r-band training data, which do not contain the deep drilling observations unlike the original Astromer training. For the r-band, this step involves transfer learning from the g-band. Hence, the retraining for r-band is about 3 times longer. We find that retraining the transformer model provides better results for the final classification task, and allows for faster training and experimentation with the classification models. We then tested the retrained Astromer for classification with different combinations of convolutional, long short-term memory (LSTM, Hochreiter & Schmidhuber, 1997), normalization, and dense layers, ranging from 2048 to 128 neurons, as well as reduction of the transomfer embeddings to mean, maximum or minimum values. We find that the best results are achieved with the following architecture: mean reduction of the embeddings, 1024, 512 and 256 neuron dense layers with the rectified linear unit activation function (ReLu, Agarap, 2018), normalization layer, and three neuron output layer. For training, we use the Adam optimizer with a 0.001 learning rate, scheduled to 0.0005 at epoch 4, and 0.0001 at epoch 9. After training the artificial neural network (ANN) classification models on g- and r-bands, we use extreme gradient boosting (XGB, Chen & Guestrin, 2016) to ensemble the deep learning classifications with features from the PS, WISE and Gaia surveys. We use the same XGB training parameters as found optimal in a similar classification task in Nakoneczny et al. (2021). We also test adding the features from other surveys as additional inputs to fully connected layers of the ANN, but find this approach to provide worse results. Also, ensembling with XGB allows separation of the ANN training from the final classification models, enabling for easier and faster comparisons with other surveys.
@@ 4.1 Light curve classification
$g<\mathrm{n_{obs}}/80+20.375$ 0.84 0.85 0.93 0.91 0.90 0.84 0.89 0.90
$\mathrm{SNR}>5\sigma\lor\mathrm{n_{obs}}>100$ 0.79 0.83 0.94 0.92 0.84 0.82 0.90 0.91
$\mathrm{n_{obs}}>100$ 0.73 0.80 0.95 0.93 0.76 0.79 0.92 0.92
@@
$\mathrm{SNR}>5\sigma\land\mathrm{n_{obs}}>100$ 0.50 0.64 0.97 0.95 0.50 0.63 0.95 0.94
Note. — ‘QSO fraction’ gives a fraction of quasars after the cuts with respect to the full training sample in the first row, ‘QSO F1 full sample’ assumes negative classification for quasars excluded by the cuts, ‘QSO F1 subset’ takes into the account only objects left after the data cut, ‘accuracy subset’ stands for the general three class accuracy of subsets. Boldface shows the highest value in each column.
Figure 5: QSO F1 scores for the XGB g-band models trained on the ZTF $\times$ PS data, as functions of median magnitude (top), number of ZTF observation epochs (middle), SDSS spectroscopic redshift (bottom). The gray dashed lines show overlaid quasar distributions. The models are trained using combinations of ZTF transformer classification, ZTF magnitude, and PS griz magnitudes and colors.
Since ZTF data span a wide range of magnitudes and observation epochs, we start by analyzing how the classification scores depend on those factors. Fig. 4 shows the dependence of ZTF g-band QSO F1 score on the number of observation epochs and magnitude. We find that given at least 50 observation epochs, the QSO F1 scores deteriorates by 10pp between the objects at magnitude limit $g\sim 20.8$ and the faintest data points at $g\sim 21.5$. We can remove the objects with lower classification quality by applying the dashed red line cut at $g<\mathrm{n_{obs}}/80+20.375$. Table 1 shows the scores for different combinations of limiting magnitude and minimum number of observation epochs. The cuts enable better F1 scores within the catalog, but taking into account the total loss of objects due to cuts, we observe a deterioration in the global QSO F1 scores. The suggested red line cut allows to achieve 93% QSO F1 score, with the global metric equal 85%. Fig. 5 shows detailed relations between QSO F1 score, magnitude, number of observation epochs and redshift for the ZTF $\times$ PS training data. We observe that the PS data improves the classification results at the faint end, mostly due to the deeper photometry provided by PS. At brighter magnitudes we observe superiority in ZTF classification over PS. We see that the Astromer’s limit of 200 observation epochs does not allow us to probe further increases in scores for more observing epochs. These tests analyze the ZTF classification outputs, but with their weakness that low numbers of objects at certain values, as shown by the gray dashed lines in the Fig. 5, can result in under-performing models due to a lack of training data with specific characteristics.
@@
4.2 Comparison with other surveys
@@
Note. — ‘QSO fraction’ provides the fraction of QSOs present in the cross-matches with respect to the full ZTF training sample, and ‘accuracy’ stands for the general three class accuracy. Boldface shows the highest value in each column.
@@
Table 3 shows the classification results of the full ZTF training sample, with features from other surveys added where available, and XGB processing the missing features where present. Adding PS, WISE, and Gaia features to the ZTF classification improves the QSO F1 scores by 7, 6, and 3 percent points, respectively. As shown in Section 4.1, in case of PS, the improvement is mostly at the faint end due to the PS fainter limiting magnitude.
@@
Fig. 10 presents feature rankings from XGB models trained on all available features in two cases: all the ZTF data with processing of missing features, and a subset of objects detected by all surveys, with no missing features present. The rankings are based on binary QSO vs non-QSO classification to focus on the features important for QSO detection. The full ZTF sample measures the importance with regard to all available QSOs, but may potentially underestimate WISE importance, which is available for 72% of QSOs. On the other hand, the cross-match between all the surveys measures the feature importance only for red and point-like objects detectable by both the WISE and Gaia surveys. The top plot shows that the transformer’s QSO probability is the most important feature. The bottom plot shows that for the cross-matched data, $W1-W2$ color is the most important feature, but the transformer classification takes second place with 15% importance. Within the given set of surveys, parallax and proper motion measurements from Gaia are not the most important features.
@@ 4.3 Redshift
4.4 The catalog
@@
$g<\mathrm{n_{obs}}/80+20.375$ 0.97 0.96 18.01 14.84 7.14 4.85 3.50 1.79
$\mathrm{SNR}>5\sigma\lor\mathrm{n_{obs}}>100$ 0.97 0.96 14.59 12.24 6.18 4.28 3.13 1.66
$\mathrm{n_{obs}}>100$ 0.97 0.96 10.96 9.46 5.00 3.53 2.63 1.44
@@
$\mathrm{SNR}>5\sigma\land\mathrm{n_{obs}}>100$ 0.98 0.98 3.83 3.64 2.26 1.71 1.38 0.90
Note. — ‘Accuracy’ stands for the general three class accuracy, and both ‘QSO F1’ and ‘accuracy’ describe the test data results for a given setup of data cuts. SNR cut is applied to the g-band only. Boldface shows the suggested setup.
According to the results from Sections 4.2 and 4.3, we publish the classification results based on the ZTF and WISE surveys, together with the ZTF $\times$ WISE redshift estimates. In the full ZTF catalog, we remove light curves which within $1\arcsec$ have at least one neighbor with more observation epochs. This step removes duplicated light curves by choosing the ones with the highest number of observation epochs. Figure 14 and Table 4 show the resulting number of objects classified as QSOs at different data and classification probability cuts. The full catalog consists of 45.93 million objects classified as QSOs, but due to the class and magnitude distribution differences between the training and inference data, it is necessary to calibrate the resulting probabilities to achieve the same F1 score as estimated in the experiments, or establish a minimum classification probability to reduce QSO contamination by stars and galaxies, at a cost of lowered QSO completeness. From Section 4.1 we conclude that the highest quality QSO classification requires a cut based on magnitude and number of observation epochs at $g<\mathrm{n_{obs}}/80+20.375$. At this cut, the estimated QSO F1 score equals 97%. Nakoneczny et al. (2019, 2021) uses a test based on Gaia parallaxes for a similar task of QSO classification based on SDSS data and suggests 90% minimum QSO classification probability. We obtain a limited catalog of higher purity by applying both the data and classification probability cuts, resulting in 4,849,574 objects classified as QSOs. We call this catalog QZO. However, the fainter sample below red line in fig 4 consists of many more true QSOs. Therefore, we also expect many more correctly classified QSOs below those limits, but with an unknown and higher contamination rate by stars and galaxies. The photometric redshifts are available for 33% of the QZO objects with matching W[1-4] bands from AllWISE.
@@ 5 Palomar spectroscopy
Table 5: List of random three SDSS galaxies and stars at ZTF g-band magnitude limit and number of observation epochs higher than 100, classified as AGN by our deep learning model.
@@
name z class AGN galaxy star light curve
J224649.18+132545.9 0.23 galaxy 45.8% 39.1% 15.1% fig. 18 b AGN; early-type galaxy with weak [N II]
J230105.96+275256.4 0.28 galaxy 48.3% 35.5% 16.2% fig. 18 e AGN; early-type galaxy with weak [N II]
@@
J002411.35-003310.9 - star 60.5% 3.5% 36.1% fig. 20 b K5 star; MgIb and NaD absorption
J155654.47+210719.0 - star 71.5% 9.7% 18.9% fig. 20 e CV; with strong Balmer emission
J185111.24+174645.7 - star 74.1% 2.5% 23.4% fig. 20 g F9 star; MgIb and NaD absorption
@@
We randomly selected six sources with conflicting SDSS and ZTF light curve classifications as a pilot program at Palomar Observatory on a night assigned for other science. The sample included three sources identified by SDSS as galaxies based on their spectra and three sources identified by SDSS as Galactic. All six were classified as AGN by our model. Table 5 presents these six sources, and they can be also identified in figures 18 and 20. All three SDSS Galactic sources were clearly Galactic based on the SDSS spectra, while one of the SDSS inactive galaxies was clearly active based on the SDSS spectrum.

### v2

@@ 2.4 Other surveys
We compare the ZTF based classification with AllWISE, Gaia EDR3, and PS DR1 by cross-matching the ZTF training datasets in g- and r-bands using a $1\arcsec$ matching distance. From PS, AllWISE, and Gaia, we use griz, W[1-4], and g, bp, rp magnitudes, respectively. We add colors from all magnitude pair combinations within each survey. Additionally, we add parallax and proper motion measurements from Gaia. We do not use error measurements for any of the features, either as inputs or weights for loss functions. Requiring availability of all the cross-matched features, we obtain the following subsets of the ZTF g-band training data: ZTF x PS 98%, ZTF x WISE 77%, ZTF x Gaia 45%. When comparing the classification based on different surveys, we apply the magnitude limits: PS griz 22.0, 21.8, 21.5, 20.9 AB, WISE W[1-2] 17.1, 15.7 Vega, Gaia g 21 Vega, respectively. Fig. 3 shows a comparison of quasar distributions between the full training sample and the cross-matched subsets. We also cross-match with AllWISE for the inference data, as justified in Sections 4.2 and 4.3. For simplicity, we use WISE to refer to the AllWISE sample. When comparing ZTF with the other surveys, we require the features from cross-matched surveys to be present. However, for the final catalog, we add AllWISE features and process the missing magnitudes. During training, XGB marks each split in the trees as more suitable for the objects with missing features, which allows to later classify such objects.
@@ 3.1 Astromer
We classify the light curves using Astromer (Donoso-Oliva et al., 2023), a deep learning transformer architecture, pre-trained on all the ZTF DR10 g-band data. The model processes time series of magnitude and observation date using attention layers trained in a semi supervised way by removing random windows from the light curves, and then recreating the missing parts. The model is limited to 200 input observation epochs, transforming each observation into a vector of 256 tokens, which results in an encoder output size of $200\times 256$. We first retrain the model in its original self-supervised training procedure, but using the training data from ZTF DR20. Then, we adapt the model for a classification task by processing the encoder output with other types of layers and additionally fine tuning the encoder weights.
@@ 3.2 Pipeline
We first retrain Astromer using its original training procedure on the g- and r-band training data, which do not contain the deep drilling observations unlike the original Astromer training. For the r-band, this step involves transfer learning from the g-band. Hence, the retraining for r-band is about 3 times longer. We find that retraining the transformer model provides better results for the final classification task, and allows for faster training and experimentation with the classification models. We then tested the retrained Astromer for classification with different combinations of convolutional, long short-term memory (LSTM, Hochreiter & Schmidhuber, 1997), normalization, and dense layers. We tested between 1 and 6 dense layers of size ranging from 2048 to 128 neurons. W also tested reduction of the transformer embeddings to mean, maximum or minimum values. We find that the best results are achieved with the following architecture: mean reduction of the embeddings, 1024, 512 and 256 neuron dense layers with the rectified linear unit activation function (ReLu, Agarap, 2018), normalization layer, and three neuron output layer. For training, we use the Adam optimizer, and we tested learning rate values lower or equal to 0.1. We arrive at the best solution with a 0.001 learning rate, scheduled to 0.0005 at epoch 4, and 0.0001 at epoch 9. After training artificial neural networks (ANN) on g- and r-bands, we use XGB to ensemble the deep learning classifications with features from the PS, WISE and Gaia surveys. We use the same XGB training parameters as found optimal in a similar classification task in Nakoneczny et al. (2021). We also test adding the features from other surveys as additional inputs to fully connected layers of the ANN, but find this approach to provide worse results. Also, ensembling with XGB allows separation of the ANN training from the final classification models, enabling for easier and faster comparisons with other surveys.
@@ 4.1 Light curve classification
$g<n_{\mathrm{obs}}/80+20.375$ 0.84 0.85 0.93 0.91 0.90 0.84 0.89 0.90
$\mathrm{SNR}>5\sigma\lor n_{\mathrm{obs}}>100$ 0.79 0.83 0.94 0.92 0.84 0.82 0.90 0.91
$n_{\mathrm{obs}}>100$ 0.73 0.80 0.95 0.93 0.76 0.79 0.92 0.92
@@
$\mathrm{SNR}>5\sigma\land n_{\mathrm{obs}}>100$ 0.50 0.64 0.97 0.95 0.50 0.63 0.95 0.94
Note. — ‘QSO fraction’ gives a fraction of quasars after the cuts with respect to the full training sample in the first row, ‘QSO F1 full sample’ assumes negative classification for quasars excluded by the cuts, ‘QSO F1 subset’ takes into the account only objects left after the data cut, ‘accuracy subset’ stands for the general three class accuracy of subsets. Class distribution affects the accuracy score, in contrast to the F1 score, which can be applied to unbalanced class problems.
Figure 5: QSO F1 scores for the XGB g-band models trained on the ZTF $\times$ PS data, as functions of median magnitude (top left), mean magnitude error (top right), number of ZTF observation epochs (bottom left), SDSS spectroscopic redshift (bottom right). The gray dashed lines show overlaid quasar distributions. The models are trained using combinations of ZTF transformer classification, ZTF magnitude, and PS griz magnitudes and colors.
Since ZTF data span a wide range of magnitudes and observation epochs, we start by analyzing how the classification scores depend on those factors. Fig. 4 shows the dependence of ZTF g-band QSO F1 score on the number of observation epochs and magnitude. We find that given at least 50 observation epochs, the QSO F1 scores deteriorates by 10pp between the objects at magnitude limit $g\sim 20.8$ and the faintest data points at $g\sim 21.5$. We can remove the objects with lower classification quality by applying the dashed red line cut at $g<n_{\mathrm{obs}}/80+20.375$. Table 1 shows the scores for different combinations of limiting magnitude and minimum number of observation epochs. The cuts enable better F1 scores within the catalog, but taking into account the total loss of objects due to cuts, we observe a deterioration in the full sample QSO F1 scores.
The suggested red line cut, in Figure 4, allows to achieve 93% QSO F1 score, with the full sample metric equal 85%. The main goal of the cut is to provide classification beyond the $5\sigma$ signal-to-noise limit. We base this cut on the results from ZTF data only, in comparison to adding WISE features as shown later in Fig. 9. We want to provide a fiducial version of the catalog with good quality light curves, useful in classification, and with a possible application in search of GW counterparts. It is possible to apply different or more complex cuts, depending on the application. We publish a file with QSOs limited to this cut, and another one with all the ZTF objects classified as quasars.
Fig. 5 shows detailed relations between QSO F1 score, magnitude, noise, number of observation epochs and redshift for the ZTF $\times$ PS training data. We observe that the PS data improves the classification results at the faint end, mostly due to the deeper photometry provided by PS. At brighter magnitudes we observe superiority in ZTF classification over PS. We see that the Astromer’s limit of 200 observation epochs does not allow us to probe further increases in scores for more observing epochs. These tests analyze the ZTF classification outputs, but with their weakness that low numbers of objects at certain values, as shown by the gray dashed lines in the Fig. 5, can result in under-performing models due to a lack of training data with specific characteristics.
@@
4.2 Other surveys
@@
Note. — ‘QSO fraction’ provides the fraction of QSOs present in the cross-matches with respect to the full ZTF training sample, and ‘accuracy’ stands for the general three class accuracy.
@@
Table 3 shows the classification results of the full ZTF training sample, with features from other surveys added where available, and XGB processing the missing features where present. Adding PS, WISE, and Gaia features to the ZTF classification improves the QSO F1 scores by 7, 6, and 3 percent points, respectively. As shown in Section 4.1, in case of PS, the improvement is mostly at the faint end due to the PS fainter limiting magnitude. The last row shows that adding ZTF data to all the other surveys increases the classification score by 2 percent points.
@@
Fig. 10 presents feature rankings from XGB models trained on all available features in two cases: all the ZTF data with processing of missing features, and a subset of objects detected by all surveys, with no missing features present. The rankings are based on binary QSO vs non-QSO classification to focus on the features important for QSO detection. The full ZTF sample measures the importance with regard to all available QSOs, but may potentially underestimate WISE importance, which is available for 72% of QSOs. On the other hand, the cross-match between all the surveys measures the feature importance only for red and point-like objects detectable by both the WISE and Gaia surveys. The top plot shows that the transformer’s QSO probability is the most important feature. The bottom plot shows that for the cross-matched data, W1W2W1-W2 color is the most important feature, but the transformer classification takes second place with 15% importance. Within the given set of surveys, parallax and proper motion measurements from Gaia are not the most important features.
@@ 4.3 Redshift
Figure 13: Feature importance for QSO photometric redshifts from XGB models trained on the ZTF g-band and WISE data. Left: colors from standalone surveys only, right: added features from a mix of ZTF and WISE surveys. Both approaches provide the same redshift estimation error.
Fig. 13 shows that all WISE bands are useful for QSO redshifts. We consider adding colors from a mix of ZTF and WISE survey, such as $g_{ZTF}-W1_{WISE}$. However, both approaches provide similar performance. Adding features which do not improve the estimates may lead to overfitting, hence, we decide not to add such colors in the final inference model. We check that requiring W[1-2] bands provides 323M object in the ZTF inference dataset, while requiring W[1-4] bands provides 262k less objects, which is 0.08% difference. Due to that, we use all WISE bands in the inference.
4.4 The QZO catalog
@@
$g<n_{\mathrm{obs}}/80+20.375$ 0.97 0.96 18.01 14.84 7.14 4.85 3.50 1.79
$\mathrm{SNR}>5\sigma\lor n_{\mathrm{obs}}>100$ 0.97 0.96 14.59 12.24 6.18 4.28 3.13 1.66
$n_{\mathrm{obs}}>100$ 0.97 0.96 10.96 9.46 5.00 3.53 2.63 1.44
@@
$\mathrm{SNR}>5\sigma\land n_{\mathrm{obs}}>100$ 0.98 0.98 3.83 3.64 2.26 1.71 1.38 0.90
Note. — ‘Accuracy’ stands for the general three class accuracy, and both ‘QSO F1’ and ‘accuracy’ describe the test data results for a given setup of data cuts. SNR cut is applied to the g-band only. Boldface shows the fiducial setup.
Based on the results from Sections 4.2 and 4.3, and taking into account computational complexity of including more surveys in the final inference, we publish the classification results based on the ZTF and WISE surveys, together with the ZTF $\times$ WISE redshift estimates. We also publish the XGB and ANN models, which can be used to create a catalog based on a different combination of surveys.
In the full ZTF catalog, we remove light curves which within $1\arcsec$ have at least one neighbor with more observation epochs. This step removes duplicated light curves by choosing the ones with the highest number of observation epochs. Figure 15 and Table 4 show the resulting number of objects classified as QSOs at different data and classification probability cuts. The full catalog consists of 45.93 million objects classified as QSOs, but due to the class and magnitude distribution differences between the training and inference data, it is necessary to calibrate the resulting classification probabilities. If such calibration is not possible, as in our case, one can establish a minimum classification probability to reduce QSO contamination by stars and galaxies, at a cost of lowered QSO completeness. From Section 4.1 we conclude that a high quality QSO classification requires a cut based on magnitude and number of observation epochs at $g<n_{\mathrm{obs}}/80+20.375$, which we consider a fiducial cut. At this cut, the estimated QSO F1 score equals 97%. Nakoneczny et al. (2019, 2021) uses a test based on Gaia parallaxes for a similar task of QSO classification based on SDSS data and suggests 90% minimum QSO classification probability. We consider this as an additional fiducial cut, while other cuts are possible depending on the application. We publish all the objects classified as quasars to allow for any postprocessing of the predictions. We obtain a limited catalog of higher purity by applying both the data and classification probability cuts, resulting in 4,849,574 objects classified as QSOs. We call this catalog QZO. We also expect many correctly classified QSOs in the fainter sample below red line in fig 4, but with an unknown and higher contamination rate by stars and galaxies, as shown in Fig. 4 and 9. The photometric redshifts are available for 33% of the QZO objects with matching W[1-4] bands from AllWISE.
@@
Figure 16:
Figure 17: Magnitude distribution based on the inference with XGB model including WISE features, with each class limited to classification probability higher than 90%, and applying the fiducial magnitude and number of observation cut. Top: counts per bin, bottom: areas normalized to unity.
Fig. 17 shows magnitude distributions for each class in the inference data, based on the XGB model with WISE features, limited to classification probability higher than 90%, and applying the fiducial magnitude and number of observation cut. We observe the model mostly classifying objects as stars, and number of quasars rising with increasing magnitude, as noticed earlier in Fig. 1. Additionally, we see that the bimodal galaxy distribution present in the training data does not bias the final inference.
Figure 18:
Figure 19: Photometric redshift distribution for a subsample of the quasar catalog with WISE features available, at $g<n_{\mathrm{obs}}/80+20.375$, as a function of minimal QSO classification probability.
Fig. 19 shows photo-z distribution for quasars with available WISE features at fiducial $g<n_{\mathrm{obs}}/80+20.375$. At lowest quasar classification probability, we observe a maximum of photo-z estimates at $z\sim 2.4$. The fiducial cut at $p_{\mathrm{QSO}}>0.9$ shows a maximum at $z\sim 1.5$, with broad wings of the distribution reaching to $z\sim 0.8$ and $z\sim 2.4$. An influx of QSOs at $z\sim 0.8$ disappears at $p_{\mathrm{QSO}}>0.998$.
@@ 5 Palomar spectroscopy
Table 5: List of random three SDSS galaxies and stars at ZTF g-band magnitude limit and number of observation epochs higher than 100, classified as AGN by our deep learning model. The $z_{\mathrm{pred}}$ column is additionally based on WISE features.
@@
name z class AGN galaxy star light curve $z_{\mathrm{pred}}$
J224649.18+132545.9 0.23 galaxy 45.8% 39.1% 15.1% fig. 23 b 0.35 AGN; early-type galaxy with weak [N II]
J230105.96+275256.4 0.28 galaxy 48.3% 35.5% 16.2% fig. 23 e 0.49 AGN; early-type galaxy with weak [N II]
@@
J002411.35-003310.9 - star 60.5% 3.5% 36.1% fig. 25 b - K5 star; MgIb and NaD absorption
J155654.47+210719.0 - star 71.5% 9.7% 18.9% fig. 25 e - CV; with strong Balmer emission
J185111.24+174645.7 - star 74.1% 2.5% 23.4% fig. 25 g - F9 star; MgIb and NaD absorption
@@
We randomly selected six sources with conflicting SDSS and ZTF light curve classifications as a pilot program at Palomar Observatory on a night assigned for other science. The sample included three sources identified by SDSS as galaxies based on their spectra and three sources identified by SDSS as Galactic. All six were classified as AGN by our model. Table 5 presents these six sources, and they can be also identified in figures 23 and 25. All three SDSS Galactic sources were clearly Galactic based on the SDSS spectra, while one of the SDSS inactive galaxies was clearly active based on the SDSS spectrum.